I am using an online study guide to prepare for the COMPASS Assessment Test, I am having difficulty with a particular word problems involving d=rt.
Two cyclist start biking from a trail's start point 3 hours apart. The second cyclist travels at 10 mph and starts 3 hours after the first cyclist who is biking at 6 mph. How much time will pass before the second cyclist catches up with the first from the time the second cyclist starts biking?
A. 2 hours
B. 4.5 hours-correct answer
C. 5 8/4 hours
D. 6 hours
E. 7.5 hours
I have no idea even where to start!

Answer by stanbon(75887)

Two cyclist start biking from a trail's start point 3 hours apart. The second cyclist travels at 10 mph and starts 3 hours after the first cyclist who is biking at 6 mph. How much time will pass before the second cyclist catches up with the first from the time the second cyclist starts biking?
-----------------
1st cyclist DATA:
rate = 6 mph ; time = x hrs ; distance = rt = 6x miles
--------------------
2nd cyclist DATA:
rate = 10 mph ; time = x-3 hrs ; distance = rt = 10(x-3)= 10x-30 miles
----------------
EQUATION:
distance = distance
6x = 10x-30
4x = 30
x = 15/2 = 7 1/2 hrs (time it takes the 2nd cyclist to catchup to the 1st)
======================
Cheers,
Stan H.
======================================
A. 2 hours
B. 4.5 hours-correct answer
C. 5 8/4 hours
D. 6 hours
E. 7.5 hours